# Multi-Model Framework for Reconstructing Gamma-Ray Burst Light Curves

### v1

@@ Abstract
Mitigating data gaps in Gamma-ray bursts (GRBs) light curves (LCs) holds immense value for its application in cosmological research because it provides more precise measurements of the parameter of interest of the two-dimensional Dainotti relation which is a relation among the end time of the plateau emission, $T_{a}$, its respective luminosity, $L_{a}$ which is calculated from the fluxes at the end of the plateau, $F_{a}$. This study extends the work done by Dainotti et al. 2023c; Manchanda et al. 2024 on the 545 GRB sample by introducing six different models: Deep Gaussian Process (DGP), Temporal Convolutional Network (TCN), Hybrid model of Convolutional Neural Network with Long Short-Term Memory (CNN-LSTM), Bayesian Neural Network (BNN), Polynomial Curve Fitting and Isotonic Regression. Our findings demonstrate that Isotonic Regression achieves the highest uncertainty reduction for all three parameters (36.3% for $\log T_{a}$, 36.1% for $\log F_{a}$, and 43.6% for $\alpha$) outperforming all the other models. The CNN-LSTM model shows consistent improvements across all GRB parameters with the lowest outlier rate for $\alpha$ (0.550%), surpassing the performance of the LSTM model in Manchanda et al. 2024. The DGP model offers reliable uncertainty reduction across all parameters and improves upon the single-layer GP baseline. These advancements are essential for using GRBs as theoretical model discriminators via the parameters of their LCs and standard candles in cosmology, investigating theoretical models, and predicting GRB redshifts through recent cutting-edge machine-learning analysis (Dainotti et al., 2024a, b; Narendra et al., 2024).
@@
$\gamma$-ray bursts— statistical methods—-machine learning— light curve reconstruction
@@ 1 Introduction
GRBs are among the most luminous transients in the universe, observed up to redshift $z=9.4$ (Cucchiara et al., 2011), making them valuable probes of the early universe. Crucial details regarding the Population III stars can also be gained from a detailed study of GRBs.
@@
GRBs have traditionally been divided into two types based on how long their gamma-ray prompt episode lasted, as measured by $T_{90}$. This measure shows how long it took to capture 90% of all background-deducted counts, starting after the first 5% of counts were found (Mazets et al., 1981; Kouveliotou et al., 1993). The mergers of compact objects are usually linked to short GRBs (SGRBs), which are characterised as $T_{90}\leq 2$s (Duncan and Thompson, 1992; Narayan et al., 1992; Usov, 1992; Thompson, 1994; Levan et al., 2008; Metzger et al., 2011; Bucciantini et al., 2012; Perna et al., 2016). On the other hand, the collapse of large stars is associated with long GRBs (LGRBs), which have $T_{90}\geq 2$s (Woosley, 1993; Paczyński, 1998; MacFadyen and Woosley, 1999; Bloom et al., 2002; Hjorth et al., 2003; Woosley and Bloom, 2006; Woosley and Heger, 2006; Kumar et al., 2008; Hjorth and Bloom, 2012; Bucciantini et al., 2008; Cano et al., 2017; Lyman et al., 2017; Perna et al., 2018; Aloy and Obergaulinger, 2021; Ahumada et al., 2021).
@@
The high energy prompt phase has been interpreted by internal shell collision or magnetic reconnection (Vestrand et al., 2005; Blake et al., 2005; Beskin et al., 2010; Gorbovskoy et al., 2012; Vestrand et al., 2014) and the long-lasting multi-wavelength afterglow phase as the interaction of the shells with the circumburst medium (Costa et al., 1997; van Paradijs et al., 1997; Piro et al., 1998; Gehrels et al., 2009). Instruments on board the Neil Gehrels Swift Observatory (Swift, Gehrels et al. 2004), BAT, XRT, and UVOT, have been crucial for rapidly identifying GRBs and enabling follow-up across X-ray to optical bands (Barthelmy et al., 2005; Burrows et al., 2005; Roming et al., 2005). Furthermore, novel features of GRB LCs have been discovered by Swift’s quick multi-wavelength afterglow follow-up (Tagliaferri et al., 2005; Nousek et al., 2006; Troja et al., 2007).
@@
Most of the X-ray LCs exhibit a plateau after the prompt phase (Zhang et al., 2006; O’Brien et al., 2006; Nousek et al., 2006; Sakamoto et al., 2007; Liang et al., 2007; Willingale et al., 2007; Dainotti et al., 2008, 2010, 2016, 2017; Dereli-Bégué et al., 2024), which can be modeled using a broken power-law (BPL) (Zhang et al., 2006, 2007; Racusin et al., 2009), a smoothly BPL, or the phenomenological model proposed by (W07, Willingale et al. 2007). Section  2.2 details the W07 and the critical parameters. The plateau phase is particularly noteworthy and is frequently explained using the framework of the magnetar model (Zhang and Mészáros, 2001; Rowlinson et al., 2014; Rea et al., 2015; Stratta et al., 2018), where precise measurements of $T_{a}$ (time at the end of plateau) are necessary to confirm this model’s validity. Moreover, a significant anti-correlation between rest-frame time $T^{*}_{X,a}$ and the corresponding X-ray luminosity $L_{X}$, known as the Dainotti 2D relation, has been identified (Dainotti et al., 2008, 2010, 2011, 2013, 2015, 2017; Tang et al., 2019; Wang et al., 2022; Zhao et al., 2019; Liang et al., 2007; Li et al., 2018).
@@
This 2D L-T relation is further developed into the 3D Dainotti relation (Dainotti et al., 2016, 2017, 2020a, 2022b), incorporating prompt luminosity $L_{X,peak}$, and aiding cosmological parameter constraints (Dainotti et al., 2023b, 2022g, 2022b; Cao et al., 2022a, b). The reduction of the uncertainties plateau parameter by 47.5% could match the cosmological precision of $\Omega_{M}$ from SNe Ia (Dainotti et al., 2020b) within 8 years (Dainotti et al., 2022g; Betoule et al., 2014), compared to the 16 years required at current observation rates (Dainotti et al., 2022g), highlighting the prospects of a practical LC reconstruction (LCR) approach.
@@
A critical barrier in utilizing GRBs for population studies or cosmology lies in the presence of observational gaps in light curves, caused by instrumental constraints or follow-up delays. These gaps hinder reliable model testing, such as the standard fireball model (Panaitescu and Kumar, 2000; Piran, 1999) or evaluating closure relations (Willingale et al., 2007; Evans et al., 2009; Racusin et al., 2009; Kumar and Duran, 2010; Srinivasaragavan et al., 2020; Dainotti et al., 2021; Ryan et al., 2020; Tak et al., 2019). LCR methods have emerged as a powerful solution to this problem.
@@
Previous works such as Dainotti et al. 2023c and Sourav et al. 2023, have introduced probabilistic and deep learning techniques, such as Gaussian Processes (GP) (Rasmussen, 2003) and Long Short-Term Memory (LSTM; Hochreiter and Schmidhuber 1997a), to reconstruct the temporal gaps in the LCs. Manchanda et al. 2024 expanded upon these approaches by evaluating ten different models for LCR. Building upon these advancements, we extend the LCR framework by introducing six more models: Deep Gaussian Process (DGP), Temporal Convolutional Network (TCN), Hybrid model of Convolutional Neural Network with Long Short-Term Memory (CNN-LSTM), Bayesian Neural Network (BNN), Polynomial Curve Fitting and Isotonic Regression.
@@
This paper is organized as follows: Section §2 offers an in-depth description of the dataset used and the different models applied to reconstruct GRB LCs. The uncertainty, performance, and outliers in Sec. §3. Sec. §4 provides the synopsis and conclusions on the observed efficacy of each model.
@@ 2.1.2 TCN
Temporal Convolutional Networks (TCNs) are used for GRB LCR because they are proficient in identifying intricate temporal trends and long-range dependencies in time-series data (Bai et al., 2018a). GRB LCs often feature sudden bursts and fluctuations across various time scales. TCNs are well-suited to handle this kind of data because they avoid issues like the weights of the model either increasing exponentially or going toward zero, namely exploding or vanishing gradients, which can occur in traditional models like Recurrent Neural Networks (RNNs). They can learn patterns in GRB data more efficiently, making them particularly useful for predicting the evolution of LCs over time.
@@ 2.1.3 CNN-BiLSTM
CNN-LSTM models integrate the advantages of convolutional neural networks and long short-term memory networks. We use this hybrid CNN-LSTM model with the expectation that it would perform better than the standard Bi-LSTM in reconstructing GRB LCs. The CNN component first detects and emphasizes important local features in the data, such as rapid bursts, while the LSTM network then models how these features evolve. This separation of feature extraction and sequence modeling can make CNN-LSTM more focused and reliable, especially in noisy LCs. In contrast, Bi-LSTM processes the entire sequence bidirectionally without explicitly identifying which input parts are most relevant, making it computationally heavier and less effective at filtering noise.
@@ 2.1.4 BNN
Bayesian Neural Networks (BNNs) are chosen for GRB LCR because they offer a way to estimate uncertainty in predictions. Since GRB data can be noisy, sparse, and unpredictable, it’s crucial to understand the confidence level of the model’s predictions. BNNs allow for uncertainty estimation by incorporating probabilistic distributions over the model’s parameters (Blundell et al., 2015). This helps in situations where data is limited or uncertain, providing a more reliable model for LCR by quantifying both the mean and variance of predictions.
@@ 2.1.6 Isotonic Regression
Isotonic regression is a non-parametric method that can fit a monotonic curve to the data, which makes it well-suited for situations where the relationship between time and intensity is not linear but still follows a consistent trend (Barlow et al., 1972). This method is flexible and can handle noisy data better than simple linear regression, making it useful for reconstructing the LCs in GRBs with flares and breaks.
@@ 2.2 Dataset and the Willingale model
This study utilizes the same dataset as described in Dainotti et al. 2023c; Manchanda et al. 2024, comprising 545 GRBs. The data originates from the Swift BAT-XRT repository (Evans et al., 2009, 2007) and has been previously presented in Srinivasaragavan et al. 2020; Dainotti et al. 2020a, 2024c, 2024b; Narendra et al. 2024. All light curves exhibit temporal gaps, with a minimum gap size of 0.03 in the log-time scale adopted for the reconstruction task.
The W07 function has been used to model LCs of GRBs, defined in Eq. 1 and introduced by Willingale et al. 2007:
$f(t)=\left\{\begin{array}[]{ll}\displaystyle{F_{i}\exp{\left(\alpha_{i}\left(1-\frac{t}{T_{i}}\right)\right)}\exp{\left(-\frac{t_{i}}{t}\right)}}{\hskip 28.45274pt\rm for}\ \ t<T_{i},\\ \\ \displaystyle{F_{i}\left(\frac{t}{T_{i}}\right)^{-\alpha_{i}}\exp{\left(-\frac{t_{i}}{t}\right)}}{\hskip 28.45274pt\rm for}\ \ t\geq T_{i}.\\ \end{array}\right.$ (1)
@@
The LCs analyzed in this study have been curated to remove the prompt emission segment due to its significant variability and the challenges in modeling it effectively.
The GRBs are classified into four distinct classes as mentioned in Dainotti et al. 2023a; Manchanda et al. 2024; Dainotti et al. 2023c:
@@
    Flares/Bumps: GRBs showing flares and bumps throughout the afterglow phase, account for 24.06% of the dataset.
@@
The parameters for each GRB LC has been derived from the work of Srinivasaragavan et al. 2020. Fig. 1 illustrates the LCs for the four identified classes and their corresponding W07 model fits. The blue points refer to the observed data that has been trimmed of the prompt emission.
Figure 1: The GRB LCs can be classified into four types based on the afterglow feature: i) Good GRBs (upper left); ii) GRB LCs that have a break in the terminal end (upper right); iii) Flares or Bumps in the afterglow (bottom left); iv) Flares or Bumps that have Double Break at the end of the LC (bottom right).References: [1] Sakamoto et al. 2008; [2] Bi et al. 2018; [3] Dainotti et al. 2022a.

### v2

@@ Abstract
Mitigating data gaps in Gamma-ray bursts (GRBs) light curves (LCs) is crucial for cosmological research, enhancing the precision of parameters, assuming perfect satellite conditions for complete LC coverage with no gaps. This analysis improves the applicability of the two-dimensional Dainotti relation, which connects the rest-frame end time of the plateau emission ($T_{a}$) and its luminosity ($L_{a}$), derived from the fluxes ($F_{a}$). The study expands on a previous 521 GRB sample by incorporating seven models: Deep Gaussian Process (DGP), Temporal Convolutional Network (TCN), Hybrid CNN with Bidirectional Long Short-Term Memory (CNN-BiLSTM), Bayesian Neural Network (BNN), Polynomial Curve Fitting, Isotonic Regression, and Quartic Smoothing Spline (QSS). Results indicate that QSS significantly reduces uncertainty across parameters—43.5% for $\log T_{a}$, 43.2% for $\log F_{a}$, and 48.3% for $\alpha$, outperforming the other models where $\alpha$ denotes the slope post-plateau based on Willingale’s 2007 functional form. The Polynomial Curve Fitting model demonstrates moderate uncertainty reduction across parameters, while CNN-BiLSTM has the lowest outlier rate for $\alpha$ at 0.77%. These models broaden the application of machine-learning techniques in GRB LC analysis, enhancing uncertainty estimation and parameter recovery, and complement traditional methods like the Attention U-Net and Multilayer Perceptron (MLP). These advancements highlight the potential of GRBs as cosmological probes, supporting their role in theoretical model discrimination via LC parameters, serving as standard candles, and facilitating GRB redshift predictions through advanced machine-learning approaches.
@@
$\gamma$-ray bursts— statistical methods—-machine learning— light curve reconstruction $\gamma$-ray bursts— statistical methods—-machine learning— light curve reconstruction
@@ 1 Introduction
GRBs are among the most luminous transients observed, reaching redshift $z=9.4$ (Cucchiara2011), making them valuable probes of the early universe. Crucial details about Population III stars can also be obtained from a detailed study of GRBs.
@@
GRBs have traditionally been divided into two types based on the duration of their gamma-ray prompt emission, as measured by $T_{90}$. This measure shows how long it took to capture 90% of all background-deducted counts, starting after the first 5% of counts (mazets1981catalog; kouveliotou1993classification). The mergers of compact objects are usually linked to short GRBs (SGRBs), which are characterised by $T_{90}\leq 2$s (duncan1992; Narayan1992; usov1992; thompson1994MNRAS.270..480T; levan2008intrinsic; metzger2011MNRAS.413.2031M; Bucciantini2012MNRAS.419.1537B; perna2016ApJ...821L..18P). On the other hand, the collapse of large stars is associated with long GRBs (LGRBs), which have $T_{90}\geq 2$s (woosley1993ApJ...405..273W; paczynski1998ApJ...494L..45P; macfayden1999ApJ...524..262M; bloom2002AJ....123.1111B; hjorth2003Natur.423..847H; Woosley2006ARA&A; woosley2006ApJ...637..914W; kumar2008Sci...321..376K; hjorth2012grb..book..169H; bucciantini2008; cano2017AdAst2017E...5C; lyman2017MNRAS.467.1795L; perna2018ApJ...859...48P; aloy2021MNRAS.500.4365A; Ahumada2021NatAs...5..917A).
@@
The high-energy prompt phase has been interpreted by internal shell collision or magnetic reconnection (Vestrand2005Natur; Blake2005Natur; Beskin2010ApJ; 2012MNRAS.421.1874G; 2014Sci...343...38V) and the long-lasting multi-wavelength afterglow phase as the interaction of the shells with the circumburst medium (costa1997; vanParadijs1997; Piro1998; Gehrels2009ARA&A). Instruments on board the Neil Gehrels Swift Observatory (Swift, Gehrels2004ApJ...611.1005G), BAT, XRT, and UVOT, have been crucial for rapidly identifying GRBs and enabling follow-up across X-ray to optical bands (barthelmy2005burst; burrows2005swift; Roming2005). Furthermore, novel features of GRB LCs have been discovered by Swift’s quick multi-wavelength afterglow follow-up (Tagliaferri2005; Nousek2006; Troja2007).
@@
Most X-ray LCs exhibit a plateau after the prompt phase (Zhang2006; OBrien2006; Nousek2006; Sakamoto2007; Liang2007; willingale2007testing; Dainotti2008; dainotti2010a; Dainotti2016; dainotti2017a; dereli2024unraveling), which can be modeled using a broken power-law (BPL) (Zhang2006; Zhang2007ApJ...655L..25Z; Racusin2009), a smoothly BPL, or the phenomenological model proposed by willingale2007testing(W07). Section  2.2 details the W07 and the critical parameters. The plateau phase is particularly notable and is frequently explained using the framework of the magnetar model (Zhang2001; 2014MNRAS.443.1779R; Rea2015; Stratta2018), where precise measurements of $T_{a}$ (time at the end of plateau) are necessary to confirm the validity of this model. Moreover, a significant anti-correlation between rest-frame time $T^{*}_{X,a}$ and the corresponding X-ray luminosity $L_{X}$, known as the Dainotti 2D relation, has been identified (Dainotti2008; dainotti2010a; dainotti2011a; dainotti2013determination; dainotti2015; dainotti2017a; Tang2019ApJS..245....1T; Wang:2021hcx; Zhao2019ApJ...883...97Z; Liang2007; Li2018b).
@@
This 2D L-T relation is further developed into the 3D Dainotti relation (Dainotti2016; dainotti2017a; dainotti2020a; 2022ApJS..261...25D), incorporating prompt luminosity $L_{X,peak}$, and aiding cosmological parameter constraints (dainotti2023b; dainotti2022g; dainotti2022b; 2022MNRAS.512..439C; 2022MNRAS.510.2928C). The reduction of the uncertainties on the plateau parameter by 47.5% could match the cosmological precision of $\Omega_{M}$ from SNe Ia (dainotti2020x) within 8 years (dainotti2022g; Betoule), compared to the 16 years required at current observation rates (dainotti2022g), highlighting the prospects of a practical LC reconstruction (LCR) approach. This enhancement can be performed assuming that the data augmentation would come from an ideal situation in which the satellite in question would be a perfect one with no gaps.
@@
A critical barrier to using GRBs for population studies or cosmology is the presence of observational gaps in LCs, due to instrumental constraints or follow-up delays. These gaps hinder reliable model testing, such as the standard fireball model (panaitescu2000analytic; piran1999gamma) or evaluating closure relations (willingale2007testing; evans2009methods; racusin2009jet; kumar2010external; srinivasaragavan2020investigation; dainotti2021closure; ryan2020gamma; tak2019closure). LCR methods have emerged as a powerful solution to this problem.
@@
Previous works, such as dainotti2023stochastic and sourav2023predicting, have introduced probabilistic and deep learning techniques, such as Gaussian Processes (GP) (GP2003gaussian) and Long Short-Term Memory (LSTM; hochreiter1997long), to reconstruct the temporal gaps in the LCs. manchanda2025gammarayburstlightcurve expanded on these approaches by evaluating ten different models for LCR. Building upon these advancements, we extend the LCR framework by introducing seven more models: Deep Gaussian Process (DGP), Temporal Convolutional Network (TCN), Hybrid model of Convolutional Neural Network with Long Short-Term Memory (CNN-BiLSTM), Bayesian Neural Network (BNN), Polynomial Curve Fitting, QSS (Quartic Smoothing Spline), and Isotonic Regression.
@@
This paper is organized as follows: Section §2 offers an in-depth description of the dataset used and the different models applied to reconstruct GRB LCs. The uncertainty, performance, and outliers in Sec. §3. Sec. §4 provides the synopsis and conclusions on the observed efficacy of each model. The appendices §5 & §6 discuss attempts to build another promising model and solutions to resolve a systematic shift problem discovered during analysis of a simulated GRB dataset, respectively.
@@ 2.1.2 TCN
Temporal Convolutional Networks (TCNs) are used for GRB LCR because they are proficient in identifying intricate temporal trends and long-range dependencies in time-series data (Bai2018). GRB LCs often feature sudden bursts and fluctuations across various time scales. TCNs are well-suited to handle this kind of data because they avoid issues like the weights of the model either increasing exponentially or going toward zero, namely exploding or vanishing gradients, which can occur in traditional models like Recurrent Neural Networks (RNNs). They can learn patterns in GRB data more efficiently, making them particularly useful for predicting the evolution of LCs over time.
@@ 2.1.3 CNN-BiLSTM
CNN-BiLSTM models integrate the advantages of convolutional neural networks and long short-term memory networks. We use this hybrid CNN-BiLSTM model with the expectation that it would perform better than the standard Bi-LSTM in reconstructing GRB LCs. The CNN component first detects and emphasizes important local features in the data, such as rapid bursts, while the LSTM network then models how these features evolve. This separation of feature extraction and sequence modeling can make CNN-BiLSTM more focused and reliable, especially in noisy LCs. In contrast, Bi-LSTM processes the entire sequence bidirectionally without explicitly identifying which input, such as flux and time parts are most relevant, making it computationally heavier and less effective at filtering noise.
@@ 2.1.4 BNN
Bayesian Neural Networks (BNNs) are chosen for GRB LCR because they offer a way to estimate uncertainty in predictions. Since GRB data can be noisy, sparse, and unpredictable, it’s crucial to understand the confidence level of the model’s predictions. BNNs allow for uncertainty estimation by incorporating probabilistic distributions over the model’s parameters (Blundell2015). This helps in situations where data is limited or uncertain, providing a more reliable model for LCR by quantifying both the mean and variance of predictions.
@@ 2.1.6 Isotonic Regression
Isotonic regression is a non-parametric method that can fit a monotonic curve to the data, which makes it well-suited for situations where the relationship between time and intensity is not linear but still follows a consistent trend (Barlow1972). This method is flexible and can handle noisy data better than simple linear regression, making it useful for reconstructing the LCs in GRBs with flares and breaks.
2.1.7 Quartic Smoothing Spline
Even though contemporary machine learning models have strong predictive powers for time-series analysis, issues with model interpretability, the requirement for large training datasets, and the potential for overfitting sparse, noisy observations can make it difficult to apply these models to scientific data, like GRB LCs (Mehta_2019). From the class of non-parametric statistical frameworks, the QSS offers a strong substitute in this situation (hastie2009elements). The QSS stands out for its mathematical transparency and structural simplicity, as it offers explicit and principled control over the bias-variance trade-off by solving a constrained optimization problem where the model’s smoothness is defined by a constraint on the sum of squared discontinuity jumps in its fifth derivative—a higher-order criterion well-suited for complex dynamics (1993csfw.book.....D). In contrast to many purely data-driven models, it naturally produces a continuously differentiable ($C^{3}$) reconstruction and imposes a strong and often physically desirable smoothness constraint (deboor2001practical). Consequently, the QSS is a robust and interpretable framework that provides a crucial benchmark for reconstructing complex temporal dynamics, particularly when the primary goal is to derive concrete insights from noisy and irregular observational data.
@@ 2.2 Dataset and the Willingale model
This study utilizes the same dataset as described in dainotti2023stochastic; manchanda2025gammarayburstlightcurve, comprising 521 GRBs. The data originates from the Swift BAT-XRT repository (Evans2009; Evans2007) and were originally compiled from: 455 GRBs observed by Swift between 2005 and 2019 (2020ApJ...903...18S; dainotti2020a; giovanna2024inferring), 4 GRBs detected by Fermi-LAT (dainotti2024optical) and 62 additional Swift GRBs from 2019 to 2023 (narendra2024grb). The final sample includes 521 unique GRBs, of which 230 have known redshift. The sample was selected to include GRBs with a well-defined plateau phase in their X-ray afterglow LCs and a sufficient number of data points (typically more than five) to allow meaningful reconstruction. This 521 GRBs dataset represents approximately 32.33% of all the GRBs detected by Swift from 2005 to 2023. All LCs exhibit temporal gaps, with a minimum gap size of 0.03 in the log-time scale adopted for the reconstruction task.
The W07 function has been used to model LCs of GRBs, defined in Eq. 1 and introduced by willingale2007testing:
[…]
@@
The LCs analyzed in this study have been processed to remove the prompt emission segment due to its significant variability and the challenges in modeling it effectively.
The 545 GRBs, of which 521 were used and classified into four distinct classes as mentioned in dainotti2023a; manchanda2025gammarayburstlightcurve; dainotti2023stochastic, are:
@@
    Flares/Bumps: GRBs showing flares and bumps throughout the afterglow phase account for 24.06% of the dataset.
@@
The parameters for each GRB LC have been derived from the work of 2020ApJ...903...18S. Fig. 1 illustrates the LCs for the four identified classes and their corresponding W07 model fits. The blue points refer to the observed data that has been trimmed of the prompt emission.
Figure 1: The GRB LCs can be classified into four types based on the afterglow feature: i) Good GRBs (upper left); ii) GRB LCs that have a break in the terminal end (upper right); iii) Flares or Bumps in the afterglow (bottom left); iv) Flares or Bumps that have Double Break at the end of the LC (bottom right).References: [1] 2008ApJS..175..179S; [2] 2018ApJ...866...97B; [3] 2022ApJ...938...41D.